OpenStudy (anonymous):

The point (2, −6) is the midpoint of a chord of the circle: x^2 − 6x + y^2 + 10y = −14. Find the length of the chord.

OpenStudy (raden):

first, u must determine the radius and the center of circle if given an equation of circle : x^2 + y^2 + Ax + By + C = 0 so, the coordinate center of circle : x=-A/2 and y=-B/2 radius (r) = sqrt (A^2/4 + B^2/4 - C)

OpenStudy (anonymous):

thanks! :)

OpenStudy (raden):

but not yet clear :)

OpenStudy (anonymous):

lol, still working on the problem! Here's hoping!

OpenStudy (raden):

i got the answer is 6sqrt(2)

OpenStudy (anonymous):

thank you so much, I'm still working on the problem! Hopefully I get this answer too! :)